Show that the linear combination of two functions of bounded variation is also of bounded variation. Is the product of two such functions also of bounded variation? 


1
Expert's answer
2021-02-24T07:41:31-0500

Let "f(x)" and "g(x)" be two functions of bounded variations.


So, the sum i.e., "f(x)+g(x)"


We can get "h(x)=f(x)+g(x)"

Using the definition of a function of continuous variation we know that the total variation is atmost the variation of "f(x)+g(x)"

So, "h(x)=f(x)+g(x)" is a function of continuous variation.



Now, let "c(x)=f(x)\cdot g(x)"

We can re-write the function as:-

"|(fg)(x)-(fg)(y)|\leq |f(x)||g(x)-g(y)|+|g(y)||f(x)-f(y)|\>\> [triangle\>inequality]"


So, this is a transposition of the previous function i.e., the sum of the functions.


So, we conclude that the product of two functions is also of bounded variation.